Abstract Submitted for the DPP96 Meeting of The American Physical Society

Sorting Category: 1.5.1 (theoretical)

Enhanced Confinement Simulations W. DORLAND, M. KOTSCHENREUTHER, Q. P. LIU, C. S. JONES, Institute for Fusion Studies, The University of Texas, M. A. BEER, G. W. HAMMETT, Princeton Plasma Physics Laboratory — Most existing tokamaks routinely achieve enhanced confinement regimes. In turn, most enhanced confinement regimes rely (to some degree) upon sheared E×B flows to stabilize the turbulence that otherwise limits the confinement to Lmode [see for example, Biglari, Diamond and Terry, Phys. Fl. B, 2 1 (1990)]. Unfortunately, it is easily shown that sheared E×B stabilization of microinstabilities such as the ITG mode does not scale favorably with machine size. Similarly, strong density peaking is technologically challenging at reactor scales. However, additional stabilization possibilities exist. Using nonlinear gyrofluid simulations in general geometry, we investigate and compare various microinstability stabilization mechanisms in several existing enhanced confinement modes. Experimental datasets from JET, DIII-D, C-Mod and TFTR are analyzed, and ITER operation is considered.

Prefer Oral Session X Prefer Poster Session

William Dorland

bdorland@zonker.ph.utexas.edu

Institute for Fusion Studies, The University of Texas

Special instructions: place next to poster by C. S. Jones

Date submitted: July 11, 1996 Electronic form version 1.1